OpenStudy (rayman5766):

what is the value y that the orderderd pair (-2,y) is a souloution to the equation 8x+ y = 32

OpenStudy (rayman5766):

@TheSmartOne

OpenStudy (rayman5766):

i told my self it 16 but me no sure

TheSmartOne (thesmartone):

Plug in x = -2 and solve for y

OpenStudy (rayman5766):

oh yea right A 48 B16 C 4.25 D- 2

OpenStudy (rayman5766):

and then what

TheSmartOne (thesmartone):

Plug in x = -2 into the equation and solve for y

OpenStudy (rayman5766):

well what is x going to be what is the slope

TheSmartOne (thesmartone):

You need to replace x with -2 and solve for y...

OpenStudy (rayman5766):

36

TheSmartOne (thesmartone):

How did you get that?

OpenStudy (rayman5766):

8-2x=

TheSmartOne (thesmartone):

8(-2) + y = 32 Can you solve for y?

OpenStudy (rayman5766):

well what would be y

TheSmartOne (thesmartone):

That's what you need to find. What is 8 * -2 = ?

OpenStudy (rayman5766):

ist like i had it then i lost it

OpenStudy (rayman5766):

and i know your not going to give me the answer and i dont want that i just want some one to explain it to me

TheSmartOne (thesmartone):

what is the value y that the orderderd pair (-2,y) is a souloution to the equation 8x+ y = 32 (-2, y) is in the form of (x, y) that means x = -2 that means you replace x in the equation with -2 to find the value of y in (-2, y) all you have to do is plug in -2 for x and solve for y It's that easy :P

OpenStudy (rayman5766):

i got it its 16 like i said '

TheSmartOne (thesmartone):

Please show your work.

OpenStudy (rayman5766):

x=y+16/4

OpenStudy (rayman5766):

8 x = 2 y + 32 8x/8 = 2 y + 32 /8

TheSmartOne (thesmartone):

The equation is 8x + y = 32

OpenStudy (rayman5766):

x = 2 y + 2^5/8

TheSmartOne (thesmartone):

I don't even know what you're doing... What is 8 * -2?

OpenStudy (rayman5766):

ohhhhhhhhhhhhhh -16

TheSmartOne (thesmartone):

-16 + y = 32 Add 16 on both sides y = ??

OpenStudy (rayman5766):

16

TheSmartOne (thesmartone):

32 + 16 = 16 is that what you're saying? Try again what is 32 + 16 = ?

OpenStudy (rayman5766):

48

TheSmartOne (thesmartone):

and that's your answer